# Weak Proximal Newton Oracles for Composite Convex Optimization

### v1

@@ 3.1.1 Symmetric subsets of $\mathbb{R}^{n}$
$\mathcal{R}({\mathbf{x}})$ is the indicator function for a convex and closed subset $\mathcal{K}\subset\mathbb{R}^{n}$ that is closed under permutation over coordinates. Furthermore, either $\mathcal{K}$ is non-negative (i.e., $\mathcal{K}\subseteq\mathbb{R}^{n}_{+}$), or $\mathcal{K}$ is closed to coordinate-wise sign-flips.
@@
Let us introduce some useful notation. For any subset of indices $T\subseteq[n]$ and ${\mathbf{x}}\in\mathbb{R}^{n}$, we denote the restriction of ${\mathbf{x}}$ to $T$ as the vector ${\mathbf{x}}_{T}\in\mathbb{R}^{|T|}$ such that for all $i\in\{1,\dots,|T|\}$, ${\mathbf{x}}_{T}(i)={\mathbf{x}}(T_{i})$, where $T_{i}$ denotes the ith element in $T$. We denote the $n\times{}|T|$ matrix ${\mathbf{I}}_{T}$ which is obtained by deleting from the $n\times n$ identity matrix all columns which are not indexed by $T$. Thus, ${\mathbf{I}}_{T}{\mathbf{x}}_{T}$ maps ${\mathbf{x}}_{T}\in\mathbb{R}^{|T|}$ to ${\mathbf{x}}\in\mathbb{R}^{n}$ by setting each entry $i\in[n]$ in ${\mathbf{x}}$ to $[{\mathbf{x}}_{T}]_{j}$ in case the jth element in $T$ is $i$, and zero in case $i\notin{}T$. Accordingly, we denote the restriction of $\mathcal{K}$ to $T$ as the set $\mathcal{K}_{T}=\{{\mathbf{x}}_{T}\in\mathbb{R}^{|T|}~{}|~{}{\mathbf{I}}_{T}{\mathbf{x}}_{T}\in\mathcal{K}\}$, and we let $\Pi_{\mathcal{K}_{T}}[\cdot]$ denote the Euclidean projection onto $\mathcal{K}_{T}$. In the following we let $\textrm{nnz}({\mathbf{x}}^{*})$ denote the number of non-zero entries in the unique optimal solution to Problem (1).
@@
    Single computation of the product of $\nabla^{2}f({\mathbf{x}}_{t})$ with a $s$-sparse vector111i.e., vector in $\mathbb{R}^{n}$ with at most $s$ non-zero entries
@@
On each iteration $i$ of Algorithm 1, given the gradient $\nabla{}Q_{t}({\mathbf{y}}_{i})$, computing the point ${\mathbf{z}}^{\prime}$ (as defined in Lemma 4) requires finding the $s$ largest entries in an $n$-dimensional array, which takes $O(n\log{}s)$ time and to to project a $s$-dimensional vector onto the corresponding $s$-dimensional restriction of $\mathcal{K}$. Since ${\mathbf{z}}^{\prime}$ is $s$-sparse, a straightforward inspection reveals that the time to compute the next gradient direction $\nabla{}Q_{t}({\mathbf{y}}_{i+1})$ is dominated by the time to multiply the Hessian $\nabla^{2}f({\mathbf{x}}_{t})$ with a $s$-sparse vector. ∎
@@
Algorithm 1 $sn$ $\tilde{\beta}/\alpha$
@@
Table 1: Comparison for solving subprobelm of minimizing $\phi_{t}({\mathbf{x}})$ when $\mathcal{R}$ is an indicator function for a $\ell_{1}$ ball in $\mathbb{R}^{n}$, when the Hessian $\nabla^{2}f({\mathbf{x}}_{t})$ is explicitly given and $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$. Universal constants and logarithmic factors are suppressed.
@@ 3.1.2 Spectrally-symmetric subsets of matrix spaces
$\mathcal{R}(\cdot)$ is the indicator function for a closed and convex subset $\mathcal{M}\subset\mathbb{R}^{m\times n}$, =$\mathcal{M}=\{{\mathbf{U}}\textrm{diag}(\sigma){\mathbf{V}}^{\top}~{}|~{}{\mathbf{U}}^{\top}{\mathbf{U}}={\mathbf{I}},~{}{\mathbf{V}}^{\top}{\mathbf{V}}={\mathbf{I}},~{}\sigma\in\mathcal{K}\}$, where $\mathcal{K}$ is a convex and compact subset of $\mathbb{R}^{\min\{m,n\}}$ that is non-negative (i.e., $\mathcal{K}\subseteq\mathbb{R}^{\min\{m,n\}}_{+}$) and closed to permutation over coordinates.
@@
Frank-Wolfe [12] $1$ $\tilde{\beta}\tau^{2}/\epsilon$
Table 2: Comparison for solving subprobelm of minimizing $\phi_{t}$ when $\mathcal{R}$ is an indicator function for a nuclear norm ball in $\mathbb{R}^{m\times n}$ of radius $\tau$ and $\textrm{rank}({\mathbf{x}}^{*})\leq s<<\min\{m,n\}$. Universal constants and logarithmic factors are suppressed. $\epsilon$ denotes the target accuracy.
@@
    Single computation of the product of $\nabla^{2}f({\mathbf{x}}_{t})$ with a rank-$s$ $m\times n$ matrix (given by it SVD).
@@ 4 Implementing WPNOs via the Decomposition-Invariant Conditional Gradient Method
Table 3: Comparison of conditional gradient methods for solving subprobelm of minimizing $\phi_{t}$ when $\mathcal{R}$ is an indicator function for a polytope satisfying Assumption 5 and $n^{*}=\dim\mathcal{F}^{*}+1<<n$.
@@
     if $\phi_{t}(({\mathbf{y}}_{i}+\lambda_{i}({\mathbf{w}}_{i}^{+}-{\mathbf{w}}_{i}^{-}))<\phi_{t}({\mathbf{y}}_{i})$ then
@@
Thus, by running Algorithm 2 for sufficiently-many iterations ($\tilde{O}(n^{*}\tilde{\beta}D^{2}/\alpha)$, where $\tilde{O}(\cdot)$ suppresses logarithmic factors), we can implement a WRPNO with parameter $C_{\mathcal{A}}=O(\beta_{2})$.
@@
$\langle{{\mathbf{x}}^{*}-{\mathbf{y}}_{i},\nabla{}Q_{t}({\mathbf{y}}_{i}}\rangle=\sum_{j=1}^{n^{*}}\Delta_{j}\langle{{\mathbf{u}}_{j}-{\mathbf{z}},\nabla{}Q_{t}({\mathbf{y}}_{i})}\rangle.$
@@
where we have used the fact that both ${\mathbf{y}}_{i},{\mathbf{y}}_{i+1}$ are guranteed to be feasible and so $\phi_{t}({\mathbf{y}}_{i+1})=Q_{t}({\mathbf{y}}_{i+1})$ and $\phi_{t}({\mathbf{y}}_{i})=Q_{t}({\mathbf{y}}_{i})$.
@@
5 Preliminary Numerical Evidence
We provide preliminary demonstration for the plausibility of our approach on a simple model of sparse logistic regression in the following $\ell_{1}$-constrained formulation:
@@ 5 Preliminary Numerical Evidence
$\min_{{\mathbf{w}}\in\mathbb{R}^{n}:~{}\|{{\mathbf{w}}}\|_{1}\leq\tau}\left\{f({\mathbf{w}}):=\sum_{i=1}^{m}\log\left({1+\exp\left({-y_{i}{\mathbf{w}}^{\top}{\mathbf{x}}_{i}}\right)}\right)\right\}.$ (28)
We draw $m=12000$ random feature vectors ${\mathbf{x}}_{i}$ in dimension $n=3000$ from a Gaussian distribution with random covariance matrix with condition number $=300$. The labels $y_{i}$ are set according to a random sparse ground-truth regressor ${\mathbf{w}}_{\sharp}$ such that $\textrm{nnz}({\mathbf{w}}_{\sharp})=k$, where $k$ varies in our experiment, and $\|{{\mathbf{w}}_{\sharp}}\|_{1}=10$. We set $\tau=1.3$ in (28) which ensures that the optimal solution to (28), ${\mathbf{w}}^{*}$, has sparsity level similar to ${\mathbf{w}}_{\sharp}$. We implement the cubic-regularized proximal Newton method, with regularization parameter 1 and with unit step-size (i.e., setting $\beta_{2}=1,\eta=1$ in (6)) in two ways. The first (the baseline) uses a standard implementation of FISTA [3] with backtracking to solve the inner optimization problem (Cubic_FISTA in Figure 1). The second uses our weak proximal oracle approach detailed in Section 3.1.1, where we choose the step-size on each inner-iteration by trying the values $0.3\cdot 10^{-j},j\in\{0,1,..,4\}$ and taking the one which most reduces the inner objective (Cubic_WPO in Figure 1). For each value of the sparsity parameter $k$ we average the results of 20 i.i.d. runs.
Figure 1 presents the (average) log approximation error ($\log(f({\mathbf{w}}_{t})-f({\mathbf{w}}^{*}))$) vs. the runtime. We can see that for low values of $k$ (highly sparse ${\mathbf{w}}^{*})$ indeed our approach is faster than the FISTA baseline due to the highly efficient Hessian - sparse vector products employed by our method (as opposed to Hessian - dense vector products in the baseline, see discussion in Section 3.1.1). However, as $k$ increases, the better dependence of the FISTA baseline on the condition number of the inner optimization problems triumphs.
(a) $k=10$
(b) $k=20$
(c) $k=30$
@@
Figure 1: Comparison of FISTA-based and first-order WPO-based implementations of the cubically-regularized proximal Newton method for various values of the sparsity parameter $k$.

### v2

@@ 3.1.1 Symmetric subsets of $\mathbb{R}^{n}$
$\mathcal{R}({\mathbf{x}})$ is the indicator function for a convex and closed subset $\mathcal{K}\subset\mathbb{R}^{n}$ that is closed under permutation over coordinates. Furthermore, either $\mathcal{K}$ is non-negative (i.e., $\mathcal{K}\subseteq\mathbb{R}^{n}_{+}$), or $\mathcal{K}$ is closed under coordinate-wise sign-flips.
@@
Let us introduce some useful notation. For any subset of indices $T\subseteq[n]$ and ${\mathbf{x}}\in\mathbb{R}^{n}$, we denote the restriction of $\mathbf{x}$} to $T$ as the vector ${\mathbf{x}}_{T}\in\mathbb{R}^{|T|}$ such that for all $i\in\{1,\dots,|T|\}$, ${\mathbf{x}}_{T}(i)={\mathbf{x}}(T_{i})$, where $T_{i}$ denotes the ith element in $T$. We denote the $n\times{}|T|$ matrix ${\mathbf{I}}_{T}$ which is obtained by deleting from the $n\times n$ identity matrix all columns which are not indexed by $T$. Thus, ${\mathbf{I}}_{T}{\mathbf{x}}_{T}$ maps ${\mathbf{x}}_{T}\in\mathbb{R}^{|T|}$ to ${\mathbf{x}}\in\mathbb{R}^{n}$ by setting each entry $i\in[n]$ in $\mathbf{x}$} to ${\mathbf{x}}_{T}]_{j}$ in case the jth element in $T$ is $i$, and zero in case $i\notin{}T$. Accordingly, we denote the restriction of $\mathcal{K}$ to $T$ as the set $\mathcal{K}_{T}=\{{\mathbf{x}}_{T}\in\mathbb{R}^{|T|}~|~{\mathbf{I}}_{T}{\mathbf{x}}_{T}\in\mathcal{K}\}$, and we let $\Pi_{\mathcal{K}_{T}}[\cdot]$ denote the Euclidean projection onto $\mathcal{K}_{T}$. In the following we let $\textrm{nnz}({\mathbf{x}}^{*})$ denote the number of non-zero entries in the unique optimal solution to Problem (1).
@@
    Single computation of the product of $\nabla^{2}f({\mathbf{x}}_{t})$ with a $s$-sparse vector222i.e., a vector in $\mathbb{R}^{n}$ with at most $s$ non-zero entries
@@
On each iteration $i$ of Algorithm 1, given the gradient $\nabla{}Q_{t}({\mathbf{y}}_{i})$, computing the point $\mathbf{z}$$^{\prime}$ (as defined in Lemma 4) requires finding the $s$ largest entries in an $n$-dimensional array, which takes $O(n\log{}s)$ time and to project a $s$-dimensional vector onto the corresponding $s$-dimensional restriction of $\mathcal{K}$. Since $\mathbf{z}$$^{\prime}$ is $s$-sparse, a straightforward inspection reveals that the time to compute the next gradient direction $\nabla{}Q_{t}({\mathbf{y}}_{i+1})$ is dominated by the time to multiply the Hessian $\nabla^{2}f({\mathbf{x}}_{t})$ with a $s$-sparse vector. ∎
@@
Algorithm 1 snsn $\tilde{\beta}/\alpha$
@@
Table 1: Comparison for solving subproblem of minimizing $\phi_{t}({\mathbf{x}})$ when $\mathcal{R}$ is an indicator function for a $\ell_{1}$ ball in $\mathbb{R}^{n}$, when the Hessian $\nabla^{2}f({\mathbf{x}}_{t})$ is explicitly given and $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$. Universal constants and logarithmic factors are suppressed.
@@ 3.1.2 Spectrally-symmetric subsets of matrix spaces
$\mathcal{R}(\cdot)$ is the indicator function for a closed and convex subset $\mathcal{M}\subset\mathbb{R}^{m\times n}$, $\mathcal{M}=\{{\mathbf{U}}\textrm{diag}(\sigma){\mathbf{V}}^{\top}~|~{\mathbf{U}}^{\top}{\mathbf{U}}={\mathbf{I}},~{\mathbf{V}}^{\top}{\mathbf{V}}={\mathbf{I}},~\sigma\in\mathcal{K}\}$, where $\mathcal{K}$ is a convex and compact subset of $\mathbb{R}^{\min\{m,n\}}$ that is non-negative (i.e., $\mathcal{K}\subseteq\mathbb{R}^{\min\{m,n\}}_{+}$) and closed under permutation over coordinates.
@@
Frank-Wolfe [14] 11 $\tilde{\beta}\tau^{2}/\epsilon$
Table 2: Comparison for solving subproblem of minimizing $\phi_{t}$ when $\mathcal{R}$ is an indicator function for a nuclear norm ball in $\mathbb{R}^{m\times n}$ of radius $\tau$ and $\textrm{rank}({\mathbf{x}}^{*})\leq s<<\min\{m,n\}$. Universal constants and logarithmic factors are suppressed. $\epsilon$ denotes the target accuracy.
@@
    Single computation of the product of $\nabla^{2}f({\mathbf{x}}_{t})$ with a rank-$s$ $m\times n$ matrix (given by its SVD).
@@ 4 Implementing WPNOs via the Decomposition-Invariant Conditional Gradient Method
Table 3: Comparison of conditional gradient methods for solving subproblem of minimizing $\phi_{t}$ when $\mathcal{R}$ is an indicator function for a polytope satisfying Assumption 5 and $n^{*}=\dim\mathcal{F}^{*}+1<<n$.
@@
  if $\phi_{t}({\mathbf{y}}_{i}+\lambda_{i}({\mathbf{w}}_{i}^{+}-{\mathbf{w}}_{i}^{-}))<\phi_{t}({\mathbf{y}}_{i})$ then
@@
[…]
@@
$\langle{{\mathbf{x}}^{*}-{\mathbf{y}}_{i},\nabla{}Q_{t}({\mathbf{y}}_{i})}\rangle=\sum_{j=1}^{n^{*}}\Delta_{j}\langle{{\mathbf{u}}_{j}-{\mathbf{z}},\nabla{}Q_{t}({\mathbf{y}}_{i})}\rangle.$
@@
where we have used the fact that both ${\mathbf{y}}_{i},{\mathbf{y}}_{i+1}$ are guaranteed to be feasible and so $\phi_{t}({\mathbf{y}}_{i+1})=Q_{t}({\mathbf{y}}_{i+1})$ and $\phi_{t}({\mathbf{y}}_{i})=Q_{t}({\mathbf{y}}_{i})$.
@@
5 Numerical Evidence
We turn to demonstrate the potential practical benefits of our theoretical results. We consider the following optimization model corresponding to the task of 1-bit matrix completion [8], which takes the form of minimizing a $\ell_{2}$-regularized logistic loss over a nuclear norm ball of matrices:
@@ 5 Numerical Evidence
$\min_{{\mathbf{x}}\in\mathbb{R}^{n_{1}\times n_{2}}:\,\|{\mathbf{x}}\|_{*}\leq\tau}\left(f({\mathbf{x}}):=\sum_{(i,j)\in\Omega}\log\bigl(1+\exp(-y_{ij}{\mathbf{x}}_{ij})\bigr)+\frac{\rho}{2}\|{{\mathbf{x}}}\|_{F}^{2}\right),$ (25)
where $\Omega\subseteq\{1,\dots,n_{1}\}\times\{1,\dots,n_{2}\}$ is the set of observed entries, and $y_{ij}\in\{-1,+1\}$ are the observed 1-bit measurements. Here $\|\cdot\|_{*}$ denotes the nuclear norm (sum of singular values) and $\tau>0$ is the radius of the feasible set. In all experiments we take $n_{1}=n_{2}=n$ and generate a rank-$r_{\sharp}$ random ground-truth matrix ${\mathbf{x}}_{\sharp}\in\mathbb{R}^{n\times n}$ with random singular values of the form $0.1+3\cdot{}U[0,1]$. The observation set $\Omega$ is chosen uniformly at random without replacement from $\{1,\dots,n\}\times\{1,\dots,n\}$ with sampling ratio $0.5$. For each $(i,j)\in\Omega$ we observe a single binary outcome $y_{ij}\in\{-1,+1\}$ generated according to the logistic 1-bit observation model:
$\mathbb{P}\bigl(y_{ij}=1\,\big|\,{{\mathbf{x}}_{\sharp}}_{ij}\bigr)=\frac{1}{1+\exp(-{{\mathbf{x}}_{\sharp}}_{ij})},\quad\mathbb{P}\bigl(y_{ij}=-1\,\big|\,{{\mathbf{x}}_{\sharp}}_{ij}\bigr)=1-\mathbb{P}\bigl(y_{ij}=1\,\big|\,{{\mathbf{x}}_{\sharp}}_{ij}\bigr).$
We set $\tau=\|{\mathbf{x}}_{\sharp}\|_{*}$ and $\rho=0.1$.
We compare two implementations of the cubically-regularized proximal Newton method with unit step-size and with regularization parameter $1$ (i.e., iteratively minimizing $\phi_{{\mathbf{x}},\eta}^{r}(\cdot)$, as defined in (6), over the constraints with $\eta=1$ and $\beta_{2}=1$). The first implementation is based on our first-order weak proximal oracle approach (Cubic_WPO in the figures below) which uses Algorithm 1 to solve the subproblems (minimizations of $\phi_{{\mathbf{x}},\eta}^{r}(\cdot)$ over the nuclear norm ball), as discussed in Section 3.1.2 (see Corollary 2 and the example following it). In particular, this implementation requires only a single SVD computation of rank $r_{\sharp}=\textrm{rank}({\mathbf{x}}_{\sharp})$ per iteration of Algorithm 1, where for the sake of demonstration we assume $r_{\sharp}$ is known. We set a fixed step-size in Algorithm 1, $\lambda_{i}=1/2$.
As a baseline, we consider a second implementation of the cubic Newton method in which the subproblems are solved via FISTA with backtracking [3] (Cubic_FISTA in the figures below). Note that FISTA requires a full-rank SVD computation in order to project onto the nuclear norm ball which is costlier than the low-rank SVD required by Algorithm 1.
For both implementations we use a maximum of $150$ iterations for solving each subproblem and we automatically stop once the distance between two consecutive iterates is at most $10^{-12}$. Additionally, in order to further demonstrate that the improved running times of our Cubic_WPO implementation are due to the ability to leverage low-rank SVDs, we also consider a third implementation (Cubic_WPO_fullSVD in the figures below), which is the same as Cubic_WPO with the sole difference that the low-rank SVDs are replaced with full-rank SVDs (as in Cubic_FISTA).
Figure 1 presents the convergence in function values for various values of $n,r_{\sharp}$, where each plot is the average of 20 i.i.d samples of the ground-truth matrix $\mathbf{x}$$_{\sharp}$. It can be clearly seen that in terms of the convergence of the Newton iterates (left panels in Figure 1), all methods have similar convergence, with Cubic_FISTA having a slight advantage. In terms of the convergence as a function of the total number of iterations performed by the inner solver (center panels in Figure 1), we can see that, as expected, Cubic_FISTA has a clear advantage which follows from the better dependence of FISTA on the condition number of the subproblems. However, when measuring the convergence in terms of the actual runtime (right panels in Figure 1), we see that our Cubic_WPO implementation has a substantial advantage. Moreover, from the slow convergence of the Cubic_WPO_fullSVD implementation we clearly see that this improved runtime is indeed due to the ability to leverage the significantly more efficient low-rank SVD computations.
(a) $n=200,r_{\sharp}=10$
(b) $n=400,r_{\sharp}=12$
(c) $n=600,r_{\sharp}=12$
@@
Figure 1: Comparison of first-order WPO-based and FISTA-based implementations of the cubically-regularized proximal Newton method. In the left and center panels the plots for Cubic_WPO and Cubic_WPO_fullSVD coincide.
@@ 6 Acknowledgement
Appendix A Proof of Lemma 3
We first restate the lemma and then prove it.
Lemma 7.
Proof.
Define the function:
We have that,
$\underset{(b)}{=}\phi_{{\mathbf{x}},\eta}({\mathbf{v}})-\phi_{{\mathbf{x}},\eta}({\mathbf{x}}^{*})-\frac{\eta}{2}\|{{\mathbf{v}}-{\mathbf{x}}}\|_{{\mathbf{x}}}^{2}+\frac{\eta}{2}\|{{\mathbf{x}}^{*}-{\mathbf{x}}}\|_{{\mathbf{x}}}^{2}$
$\underset{(c)}{\leq}\eta^{2}C_{\mathcal{A}}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|^{3}$
Note that,
Plugging into the previous inequality, we have that
Combining Eq. (26) and Eq. (A) we obtain,